Please explain and show the correct answer of number 24 on 10.5
Write the equation of the hyperbola in standard form...
x2/9 - 4y2/9 = 9
Please, please explain

Answer by stanbon(75887)

Write the equation of the hyperbola in standard form...
x^2/9 - 4y^2/9 = 9
----
Multiply thru by (1/9) to get:
x^2 - 4y^2 = 1
--------------
Modify:
x^2/1 - y^2/(1/4) = 1
==========================
Cheers,
Stan H.